# MultiSessionCollab: Learning User Preferences with Memory to Improve Long-Term Collaboration

### 03 Mar 2026 v2

@@ Abstract
As conversational agents accumulate experience collaborating with users, adapting to user preferences is essential for fostering long-term relationships and improving collaboration quality over time. We introduce MultiSessionCollab, a benchmark that evaluates how well agents can learn user preferences and leverage them to improve collaboration quality throughout multiple sessions. To develop agents that succeed in this setting, we present long-term collaborative agents equipped with a memory that persists and refines user preference as interaction experience accumulates. Moreover, we demonstrate that learning signals can be derived from user simulator behavior in MultiSessionCollab to train agents to generate more comprehensive reflections and update their memory more effectively. Extensive experiments show that equipping agents with memory improves long-term collaboration, yielding higher task success rates, more efficient interactions, and reduced user effort. Finally, we conduct a human user study that demonstrates that memory helps improve user experience in real-world settings.
@@ 1 Introduction
Recent works have introduced memory mechanisms to overcome the finite context window of Large Language Models (LLMs) and enable agents to dynamically store and retrieve user-specific information (Chhikara et al., 2025; Wang and Chen, 2025; Yan et al., 2025). Evaluations of these systems generally focus on how accurately agents can recall information and answer questions about past interactions (Xu et al., 2022a; Maharana et al., 2024; Wu et al., 2025a). While such recall abilities are a necessary prerequisite, they do not capture all abilities needed for the intended purpose of these systems, which is to improve long-term interaction with users. In particular, it remains unclear whether agents can recognize what information is valuable for future interactions and leverage it effectively.
@@
To address this gap, we introduce MultiSessionCollab, a benchmark designed to evaluate how well conversational agents can learn user preference information and leverage it to meaningfully improve interactions during multi-session collaboration. The benchmark features user simulators that collaborate with the agent on problem-solving tasks across multiple sessions. Each user has distinct personas and interaction preferences that are derived from psychology, cognitive science, and human-computer interaction research, reflecting realistic human behaviors in collaborative settings (see taxonomy in Appendix A) (Wood and Neal, 2007; Sweller, 2006; Schwartz et al., 2002). The benchmark is designed so that users progress most effectively when agent responses align with their preferences, incentivizing agents to continuously learn and adhere to preferences. An overview of the benchmark is presented in Figure 1.
@@
Unlike prior memory systems that optimize for information recall, we develop more user-centric agents that are equipped with memory to learn user preferences and improve collaboration quality over time. As shown in Figure 1, agents reflect on interactions to identify user preference information that will be valuable for future interactions and update their memory accordingly. During subsequent sessions, the full memory is provided to the agent, and relevant parts are dynamically retrieved throughout the interaction to guide agent behavior towards being more preference-aligned.
@@
We also present a reinforcement learning (RL) framework that derives learning signals from user simulator behavior in MultiSessionCollab to train agents to update memory more effectively. We use Group Relative Policy Optimization (GRPO) (Shao et al., 2024) with an LLM-judge that rewards reflections for comprehensively capturing user preferences revealed during interaction.
@@
Extensive experiments across several models and five problem-solving tasks demonstrate that our memory enables agents to continuously learn user interaction preferences and improve collaboration quality, yielding higher task success rates, more efficient interactions, and reduced user effort. Analyses of performance across sessions reveal continued improvement throughout the sessions, with the steepest gains occurring in early sessions before gradually stabilizing. Notably, agents that learn preferences through interaction are competitive with those given direct access to ground-truth user preferences, showing how memory captures richer information about preferences.
@@
  • Through extensive experiments and a human user study, we demonstrate that memory has a positive impact on multi-session collaboration.
@@ 2 Related Work
Early works that evaluate LLMs in multi-session interactions focus on how well models can generate responses that are consistent with past interactions (Xu et al., 2022a, b). More recently, benchmarks have shifted towards question-answering style evaluations that assess how well LLMs can remember information from past interactions (Packer et al., 2023; Maharana et al., 2024; Wu et al., 2025a; Hu et al., 2025; Jiang et al., 2025; Zhao et al., 2025). In contrast, our work evaluates memory for improving user interactions, incorporating abilities such as identifying useful information and actually leveraging it to improve collaboration.
@@ 4.1 Long-Term Collaborative Agents
To build such agents, we equip them with memory that persists and refines user preference information across sessions. It is illustrated in Figure 1 and all prompts are provided in Appendix B. The memory architecture comprises two components:
@@
    Session-Level Reflection: the agent analyzes the interaction and identifies preference information that would be useful for future interactions, which can include the specific preferences themselves, the contexts in which they apply, and details about which actions or approaches satisfy each preference. The agent then updates their existing memory with this new information.
@@
    Persistent Memory: the memory is provided at the start of each session as part of the agent’s system prompt, enabling them to adapt their behavior according to learned preferences without requiring users to repeatedly specify them. Additionally, at each turn, we use an LLM to analyze the conversation history and retrieve parts of the memory that are directly relevant to the conversational context and provide them to the agent.
@@ 4.2 Reinforcement Learning in MultiSessionCollab
A session-level reflection is considered successful if it comprehensively captures the user preferences revealed during the interaction and provides actionable guidance for future sessions. $\varepsilon$ specifies all information that a reflection should capture. We therefore use an LLM judge to evaluate $r$ covers all preferences in $\varepsilon$ without any hallucinations, and get an objective scalar reward $R_{coverage}(r,\varepsilon)$. We also include a format reward $R_{format}(r)$ that encourages well-structured outputs containing both a reasoning trace and the reflection. The total reward is:
@@
LLM judge prompts are provided in Appendix B. Using this reward, we employ GRPO to train a policy $\pi_{\theta}$ to generate more comprehensive session-level reflections that have no hallucinations. As a result, the agent is able to make more meaningful memory updates that helps them be more preference-aligned during future interactions.
@@ 6 Results
Table 2: Comparison between our memory architecture and Mem0 on MultiSessionCollab. We report task success (TS), user effort (UE), and conversation length (Len) averaged across all tasks. Subscripts denote change relative to the baseline without memory.
@@
Table 1 presents our experimental results on the MultiSessionCollab benchmark. Equipping conversational agents with memory enables them to learn preferences as they accumulate experience and proactively adhere to them without requiring users to repeatedly enforce them, reducing friction and allowing users to focus on the task. This generally results in improvements in collaboration quality: interactions become more efficient, require less user effort, and yield higher task success rates.
We observe even further improvements when using our RL framework to train to generate session-level reflections that effectively capture the preferences revealed during interactions for Qwen-2.5-7B-Instruct and Llama-3.1-8B-Instruct. After training, both models are able to leverage the memory more effectively, with notable improvements across all metrics. In particular, while Qwen-2.5-7B-Instruct initially exhibited a $1.03\%$ decrease in task success when equipped with memory, after training, the same model achieves a $3.43\%$ improvement in task success. This demonstrates that the memory architecture’s effectiveness depends largely on reflection quality. Additionally, these results suggest that the way our RL framework uses learning signals derived from user simulator behavior offers a promising direction for improving agent memory management abilities.
@@
Unlike existing memory architectures, which are primarily designed for information retrieval and question-answering, ours is designed to improve interactions with users. To demonstrate this, we compare against Mem0, a memory architecture that stores information in a vector database after each user-agent turn and dynamically retrieves them prior to response generation (Chhikara et al., 2025). We equip Llama-3.3-70B-Instruct with each memory architecture and present results averaged across all tasks in Table 2.
@@
Next, we analyze our agent’s performance as they accumulate interaction experience throughout sessions. For each session $i\in\{1,...,20\}$, we compute the agent’s average performance across all users and tasks for each metric. Then, we calculate the deltas $\Delta_{i}^{TS}$, $\Delta_{i}^{UE}$, and $\Delta_{i}^{Eff}$, where each delta denotes the difference between the agent with memory and the agent without memory at session $i$ for the respective metric. These deltas isolate the impact of memory at each session. Figure 3 plots these deltas across sessions for Llama-3.3-70B-Instruct, with additional plots for other models presented in Appendix E. Increasing $\Delta_{i}^{TS}$, and decreasing $\Delta_{i}^{UE}$ and $\Delta_{i}^{Eff}$ indicate that the agent is learning to collaboration more effectively.
@@
In Appendix C, we report further ablations that help us better understand the memory architecture. One provides agents with direct access to user preferences, same as what is provided to the user simulator. This was intended to be an oracle to establish an upper bound on performance. Surprisingly, agents equipped with memory achieve competitive performance with this oracle, and sometimes outperform it, despite the fact that agents with memory begin with no user preference information. Our qualitative analysis indicates that memory helps capture richer preference information from interactions, such as the contexts in which the preferences apply and specific strategies for adhering to them. Moreover, this information is continually refined across sessions as agents accumulate interaction experience. These results demonstrate the effectiveness of using memory to learn through interaction.
@@ 7 Real-World User Study
To complement our main experiments, we conduct a user study with 19 participants to understand the impact of memory in more realistic collaboration settings. The study lasted approximately 1.5 hours per participant, and was declared exempt after being reviewed by our Institutional Review Board (IRB). The complete details of our study design and results are provided in Appendix F.
@@ 8 Conclusion
We introduce MultiSessionCollab, a benchmark for evaluating conversational agents’ ability to learn user preferences and leverage them to meaningfully improve collaboration across multiple sessions. We develop long-term collaborative agents equipped with memory that learn preferences to improve interaction quality over time. We further demonstrate that MultiSessionCollab can serve as an effective RL environment, where we use learning signals derived from user behavior to encourage more comprehensive reflections, leading to more meaningful memory updates. Extensive experiments and user studies show that equipping agents with memory improves collaboration quality across sessions, yielding higher task success rates, more efficient interactions, and reduced user effort.
@@ Discussion.
The results from our ablation studies are presented in Table 4:
@@
The transition from S1 to S2 reveals a notable drop in performance when moving from direct problem-solving to multi-turn interaction. This result aligns with findings from prior works (Wu et al., 2025b; Laban et al., 2025; Zhou et al., 2025a; Mehri et al., 2025), and demonstrates how LLMs struggle with challenges in multi-turn interactions such as handling underspecified problem descriptions.
@@ Appendix D Training Hyperparameters
Appendix E Performance Across Sessions
@@
Appendix F Real-World User Study
@@
Below are example screenshots of our user study interface:

### 15 Mar 2026 v3

@@ Abstract
As conversational agents accumulate experience collaborating with users, adapting to user preferences is essential for fostering long-term relationships and improving collaboration quality over time. We introduce MultiSessionCollab, a benchmark that evaluates how well agents can learn user preferences and leverage them to improve collaboration quality throughout multiple sessions. To develop agents that succeed in this setting, we present long-term collaborative agents equipped with a memory that is specifically designed to learn user preferences across sessions and improve interactions. Moreover, we demonstrate that learning signals can be derived from user simulator behavior in MultiSessionCollab to train agents to generate more comprehensive reflections and update their memory more effectively. Extensive experiments show that equipping agents with our memory improves collaboration over time, yielding higher task success rates, more efficient interactions, and reduced user effort. Finally, we conduct a human user study that demonstrates that memory helps improve user experience in real-world settings.
@@ 1 Introduction
Recent works have introduced memory mechanisms to overcome the finite context window of Large Language Models (LLMs) and enable agents to dynamically store and retrieve user-specific information (Chhikara et al., 2025; Wang and Chen, 2025; Yan et al., 2025). Evaluations of these systems generally focus on how accurately agents can recall information and answer questions about past interactions (Xu et al., 2022a; Maharana et al., 2024; Wu et al., 2025a), rather than their downstream impact on long-term interactions with users. While recall is a necessary prerequisite, it does not capture whether agents can recognize what information is useful and leverage it to meaningfully improve future interactions.
@@
To address this gap, we introduce MultiSessionCollab, a benchmark designed to evaluate how well conversational agents can learn user preference information and leverage it to meaningfully improve interactions during multi-session collaboration. It features user simulators that collaborate with the agent on problem-solving tasks across multiple sessions. Each user has distinct personas and interaction preferences that are derived from psychology, cognitive science, and human-computer interaction research, reflecting realistic human behaviors in collaborative settings (see taxonomy in Appendix A) (Wood and Neal, 2007; Sweller, 2006; Schwartz et al., 2002). The benchmark is designed so that users progress most effectively when agent responses align with their preferences, incentivizing agents to continuously learn and adhere to preferences. We present an overview in in Figure 1.
@@
Unlike prior memory systems that optimize for information recall, we develop agents that are equipped with a more user-centric memory that learn user preferences to improve collaboration quality over time. As shown in Figure 1, agents reflect on interactions to identify user preference information that will be valuable for future interactions and update their memory accordingly. During subsequent sessions, the full memory is provided to the agent, and relevant parts are dynamically retrieved throughout the interaction to guide agent behavior towards being more preference-aligned.
@@
We also present a reinforcement learning (RL) framework that trains agents with learning signals from user simulator behavior in MultiSessionCollab to update memory more effectively. We use Group Relative Policy Optimization (GRPO) (Shao et al., 2024) with an LLM-judge that rewards reflections for comprehensively capturing user preferences revealed during interaction.
@@
Extensive experiments across several models and five problem-solving tasks demonstrate that our memory architecture enables agents to continuously learn user interaction preferences and improve collaboration quality, yielding higher task success rates, more efficient interactions, and reduced user effort. Analyses of performance across sessions reveal continued improvement throughout the sessions, with the steepest gains occurring in early sessions before gradually stabilizing. Moreover, we show that our memory outperforms existing existing memory architectures at improving collaboration across sessions, and is even competitive with an oracle given direct access to ground-truth user preferences, highlighting the value of designing memory to be user-centric and learning through user interactions.
@@
  • Through extensive experiments and a human user study, we demonstrate that our memory improves multi-session collaboration.
@@ 2 Related Work
Early works that evaluate LLMs in multi-session interactions focus on how well models can generate responses that are consistent with past interactions (Xu et al., 2022a, b) and accurately update memory over time (Bae et al., 2022). More recently, benchmarks have shifted towards question-answering style evaluations that assess how well LLMs can remember information from past interactions (Packer et al., 2023; Maharana et al., 2024; Wu et al., 2025a; Hu et al., 2025; Jiang et al., 2025; Zhao et al., 2025). In contrast, our work evaluates the downstream impact of memory on collaborative problem solving, incorporating abilities such as identifying useful information and actually leveraging them to improve user interactions.
@@ 4.1 Long-Term Collaborative Agents
To build such agents, we equip them with a memory architecture that is specifically designed to persist and refine user preference information across sessions and improve interactions. It is illustrated in Figure 1 and all prompts are provided in Appendix B. The memory architecture comprises two components:
@@
    Session-Level Reflection: the agent analyzes the interaction and identifies preference information that would be useful for improving future interactions, which can include the specific preferences themselves, the contexts in which they apply, and details about which actions or approaches satisfy each preference. The agent then updates their existing memory with this new information.
@@
    Persistent Memory: the memory is provided at the start of each session as part of the agent’s system prompt, enabling them to adapt their behavior according to learned preferences without requiring users to repeatedly specify them. Additionally, at each turn, we use an LLM to analyze the conversation and retrieve specific parts of the memory that are directly relevant to the conversational context and provide them to the agent.
@@ 4.2 Reinforcement Learning in MultiSessionCollab
A session-level reflection is considered successful if it comprehensively captures the user preferences revealed during the interaction and provides actionable guidance for future sessions. $\varepsilon$ specifies all information that a reflection should capture. We therefore use an LLM judge to evaluate how well $r$ covers all preferences in $\varepsilon$ without any hallucinations, and get an objective scalar reward $R_{coverage}(r,\varepsilon)$. We also include a format reward $R_{format}(r)$ that encourages well-structured outputs containing both a reasoning trace and the reflection. The total reward is:
@@
The LLM judge prompts are provided in Appendix B. Using this reward, we employ GRPO to train a policy $\pi_{\theta}$ to generate more comprehensive session-level reflections that have no hallucinations. As a result, the agent is able to make more meaningful memory updates that helps them be more preference-aligned during future interactions.
@@ 6 Results
Table 2: Comparison between our memory architecture and Mem0 (Chhikara et al., 2025) on MultiSessionCollab. We report task success (TS), user effort (UE), and conversation length (Len) averaged across all tasks. Subscripts denote change over having no memory.
@@
Table 1 presents our experimental results on the MultiSessionCollab benchmark. Equipping conversational agents with our memory architecture enables them to learn preferences as they accumulate experience and proactively adhere to them without requiring users to repeatedly enforce them, reducing friction and allowing users to focus on the task. This results in improvements in collaboration quality: interactions become more efficient, require less user effort, and yield higher task success rates.
We observe even further improvements when using our RL framework to train agents to generate session-level reflections that effectively capture the preferences revealed during interactions. We train Qwen-2.5-7B-Instruct and Llama-3.1-8B-Instruct, which leads to both models being able to leverage memory more effectively and notable improvements across all metrics. In particular, while Qwen-2.5-7B-Instruct initially exhibited a $1.03\%$ decrease in task success when equipped with memory, after training, the same model achieves a $3.43\%$ improvement in task success. This demonstrates that our memory architecture’s effectiveness depends largely on reflection quality. Overall, these results suggest that RL with learning signals derived from user simulator behavior offers a promising direction for designing memory agents for improving user interactions.
@@
Unlike existing memory architectures, which are primarily designed for information retrieval and question-answering, ours is designed to improve interactions with users. To demonstrate this, we compare against Mem0, a state-of-the-art memory architecture that stores information in a vector database after each user-agent turn and dynamically retrieves them prior to response generation (Chhikara et al., 2025). We equip Llama-3.3-70B-Instruct with each memory architecture and present results averaged across all tasks in Table 2.
@@
We additionally verify in Appendix E that providing the full multi-session conversation history as context is not a viable alternative to memory.
@@
Next, we analyze our agent’s performance as they accumulate interaction experience throughout sessions. For each session $i\in\{1,...,20\}$, we compute the agent’s average performance across all users and tasks for each metric. Then, we calculate the deltas $\Delta_{i}^{TS}$, $\Delta_{i}^{UE}$, and $\Delta_{i}^{Eff}$, where each delta denotes the difference between the agent with memory and the agent without memory at session $i$ for the respective metric. These deltas isolate the impact of memory at each session. Figure 3 plots these deltas across sessions for Llama-3.3-70B-Instruct, with additional plots for other models presented in Appendix F. Increasing $\Delta_{i}^{TS}$, and decreasing $\Delta_{i}^{UE}$ and $\Delta_{i}^{Eff}$ indicate that the agent is learning to collaboration more effectively.
@@
We conduct an ablation study where agents are provided with direct access to user preferences, the same as what is given to the user simulator. This was originally intended to be an oracle to establish an upper bound on performance. Surprisingly, agents equipped with our memory architecture achieve competitive performance with this oracle, and sometimes outperform it, despite beginning with no information about the user. Our qualitative analysis indicates that our memory helps capture richer preference information from interactions, such as the contexts in which the preferences apply and specific strategies for satisfying them. Moreover, this information is continually refined across sessions as agents accumulate interaction experience. These results demonstrate the effectiveness of using memory to learn through actual user interaction. We report this and several additional ablations that help us better understand the memory architecture in Appendix C.
@@ 7 Real-World User Study
To complement our main experiments, we conduct a user study with 19 participants to understand the impact of memory in more realistic collaboration settings. The study lasted approximately 1.5 hours per participant, and was declared exempt by our Institutional Review Board (IRB). A comprehensive description of our study design and results are provided in Appendix G.
@@ 8 Conclusion
We introduce MultiSessionCollab, a benchmark for evaluating conversational agents’ ability to learn user preferences and leverage them to meaningfully improve collaboration across multiple sessions. We develop long-term collaborative agents equipped with memory designed to learn preferences to improve interaction quality over time. We further demonstrate that MultiSessionCollab can serve as an effective RL environment, where we use learning signals derived from user behavior to encourage more comprehensive reflections, leading to more meaningful memory updates. Extensive experiments and user studies show that equipping agents with our memory improves collaboration quality across sessions, yielding higher task success rates, more efficient interactions, and reduced user effort.
@@ Discussion.
The results from our ablation studies are presented in Table 4. The transition from S1 to S2 reveals a notable drop in performance when moving from direct problem-solving to multi-turn interaction. This result aligns with findings from prior works (Wu et al., 2025b; Laban et al., 2025; Zhou et al., 2025a; Mehri et al., 2025), and demonstrates how LLMs struggle with challenges in multi-turn interactions such as handling underspecified problem descriptions.
@@ Appendix D Training Hyperparameters
Appendix E Memory vs. Full Conversation History
LLMs struggle to effectively handle large contexts, which often leads to degraded performance and increased computational costs (Shi et al., 2023; Liu et al., 2024). To address this, memory mechanisms have been introduced that allow LLMs to store and retrieve relevant information during interactions, rather than conditioning on the full context (Shinn et al., 2023; Packer et al., 2023; Zhong et al., 2024; Suzgun et al., 2025; Wang et al., 2025c; Ho et al., 2025; Chhikara et al., 2025). Table 6 compares the average token length of the multi-session conversation history against the corresponding memory accumulated across all sessions. On average, the conversation history contains ${\sim}21,700$ tokens compared to ${\sim}216$ for memory, showing that the memory is on average 100x shorter.
Task #Users #Convs/User MSC Tokens Mem Tokens
LogiQA 100 20 24,074 215
MATH-500 100 20 19,007 213
MATH-Hard 100 20 24,338 210
MedQA 100 20 22,780 243
MMLU 100 20 18,539 196
Average 100 20 21,748 216
Table 6: Average token length of the full multi-session conversation (MSC) history and the corresponding memory accumulated across all 20 sessions per user.
We compare using our memory architecture against providing the full multi-session conversation history as context to the agent. We test two variants: providing the conversation history from (1) the last 3 sessions and (2) all previous sessions. We evaluate Llama-3.3-70B-Instruct, using the experimental setup described in 5. We note that both conversation history variants are significantly more expensive than using memory. For instance, a single experiment providing the all previous sessions took roughly 17 hours on 8xH100s.
MedQA LogiQA
TS (%) $\uparrow$ UE $\downarrow$ Len $\downarrow$ TS (%) $\uparrow$ UE $\downarrow$ Len $\downarrow$
Llama-3.3-70B 45.84 2.91 16.38 26.69 2.96 17.08
+ 3 past convs 43.28↓2.56 2.22↓0.69 16.04↓0.34 26.40↓0.29 1.68↓1.28 15.73↓1.35
+ full history 44.05↓1.79 2.22↓0.69 15.75↓0.63
+ memory (ours) 49.95↑4.11 1.92↓0.99 15.35↓1.03 30.25↑3.56 1.89↓1.07 15.85↓1.23
Table 7: Memory vs. full conversation history. Comparison of our memory architecture against providing the conversation history as context to Llama-3.3-70B-Instruct in MultiSessionCollab. We report task success (TS), user effort (UE), and conversation length (Len). Subscripts denote change relative to the baseline.
Results are presented in Table 7. Providing conversation history as context leads to a drop in task success, with minor decreases in user effort and conversation length. Our memory architecture outperforms both conversation history variants across all metrics. Upon closer inspection, we find that agents provided with the full history exhibit degraded behavior. They are often unable to engage in coherent conversation, and generate meaningless responses that do not address user requests. This is consistent with prior works showing that LLMs struggle to effectively handle long contexts, leading to both degraded performance and increased computational costs (Shi et al., 2023; Liu et al., 2024).
Appendix F Performance Across Sessions
@@
Appendix G Real-World User Study
@@
Figure 7, 8, 9 provide example screenshots of our user study interface.